OpenStudy (anonymous):

Find two numbers whose sum is 42 and whose difference is 16.

OpenStudy (chriss):

x+y=42 x-y=16

OpenStudy (anonymous):

So Wats The answer?

OpenStudy (chriss):

You have to solve this system of equations to get the answer. One way you could go about that would be to add the systems together. This will eliminate the y and turn it into a single variable algebra problem. x+y=42 x-y=16 ------- 2x = 58 Once you solve that for x, you can plug that result back in to either of the first two equations and solve for y.

OpenStudy (anonymous):

would the answer be 13 and 29??

OpenStudy (chriss):

That is correct :)

OpenStudy (anonymous):

Thanx alot ;p

OpenStudy (chriss):

you're welcome

OpenStudy (anonymous):

Hey Can u help me in this problem?

OpenStudy (anonymous):

Solve the system for the x-coordinate of the point of intersection. 3x + y = -2 2x - 5y = 4

OpenStudy (chriss):

This one you will want to use a similar method to the one above. You want to eliminate the y. The difference here is that just adding or subtracting won't be enough to do that for you. The trick that you can use to do that though is to first multiply the first equation by 5, then your y-terms will be opposites and when you add the equations together, they will cancel each other out. 15x+5y=-10 2x-5y=4 ------------ 17x = -6 \[x=-\frac{6}{17}\]